# Internal structure of the $T_{cc}(3875)^{+}$ from its light-quark mass dependence

### v1

@@ 1 Introduction
In 2021, the LHCb experiment discovered the first exotic doubly-charmed narrow resonance, denoted as $T_{cc}(3875)^{+}$, whose minimal quark composition is $cc\bar{u}\bar{d}$ [9, 10]. This discovery revealed a state with a mass just a few hundred keV below the $D^{*+}D^{0}$ threshold, with a width primarily dictated by its strong decay mode to $DD\pi$. The properties of this state have been investigated employing low-energy effective field theories (EFT) [11, 12, 13, 14, 15, 16] and phenomenological models, see, e.g., [7] and references therein. In particular, the pole position of the $T_{cc}(3875)^{+}$ and the $DD^{*}$ scattering parameters were extracted in Ref. [13] using the leading order chiral EFT approximation by performing a coupled-channel analysis of the experimental line shape in the $D^{0}D^{0}\pi^{+}$ final state. In that work, a special attention was paid to the inclusion of three-body cuts, which were found to be crucial for the accurate determination of the $T_{cc}$ pole position in the complex energy plane. We emphasize that for the $T_{cc}$, including the one-pion exchange (OPE) to all orders with a proper treatment of its cuts is needed for theoretical consistency. To illustrate this, we note that the $D^{*}$ self-energy induces an imaginary part to the $D^{*}D$ propagator, which corresponds to the $DD\pi$ intermediate state going on shell. This on-shell intermediate state contains a pair of identical mesons ($DD$) in isospin 1, which has to be in an even partial wave to comply with Bose symmetry. Proper symmetrization of the $DD$ state, that ensures this, generates a one-pion exchange. Moreover, since the $D^{*}$ self-energy is resummed in the $D^{*}D$ propagator, consistency with Bose symmetry requires resumming the OPE as well. For a similar discussion on the Pauli principle in the context of two identical nucleons in the three-body intermediate state, see Refs. [17, 18, 19].
@@
In this work, we employ the same chiral EFT-based approach to scrutinize the analytic structure of the scattering amplitude as a function of the pion mass. Given the intricate interplay between the right-hand (three-body) and left-hand cuts, which depend sensitively upon the pion mass, we incorporate all energy and/or momentum scales relevant for this complicated dynamics to predict the pole trajectory of the $T_{cc}$ state for pion masses between the physical value $m_{\pi}^{\rm ph}$ and $3m_{\pi}^{\rm ph}$. This information can serve as a benchmark for future lattice QCD calculations and is important for obtaining additional insights into the structure of $T_{cc}$ state [43]. Our analysis also yields predictions for the $DD^{*}$ phase shifts at any value of $m_{\pi}$ within the considered range, in spite of the fact that the ERE has a very limited range of validity. The lhc also necessitates an improvement of the Weinberg approach to properly describe the compositeness of a hadronic state. Besides the fact that the zeros in the $T$-matrix emerging from the interplay of the repulsive OPE potential and attractive short-range physics may invalidate the original Weinberg formalism, as discussed in Refs. [44, 45, 46], already the small scale introduced into the system by the close by lhc calls for a refined formulation of the compositeness criterion.
@@ 2 Framework
$V_{\text{OPE}}(E,{\bm{p}},{\bm{p}}^{\prime})=-\frac{g^{2}}{8f_{\pi}^{2}}\frac{({\bm{q}\cdot\bm{\epsilon}})({\bm{q}\cdot\bm{\epsilon}^{\prime*}})}{2\omega_{\pi}({\bm{q}}^{2})}\,D^{\pi}(E,{\bm{p}},{\bm{p}}^{\prime})$ (3)
@@
The partial-wave-projected potentials $V_{\alpha\beta}(E,p,p^{\prime})$ are obtained along the lines of Refs. [53, 54] as follows
@@
$+\int_{0}^{\Lambda}\frac{\text{d}{q}\,q^{2}}{2\pi^{2}}V_{\alpha\gamma}(E,k,q)G(E,q)T_{\gamma\beta}(E,q,k^{\prime}).$
@@
$\Gamma(E,q)=\frac{g^{2}M_{D}}{8\pi f_{\pi}^{2}M_{D^{*}}}\Big{[}\Sigma(s)-\Sigma_{0}(s)\theta(M_{D}+m_{\pi}-M_{D^{*}})\Big{]},$
with $s=[E-M_{D}-q^{2}/(2\mu)]^{2}$, and
@@
where $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2bc-2ca$ is the Källén triangle function. Here
@@
where the first and second terms renormalize the $D^{*}$ mass and wave function, respectively, if $M_{D^{*}}<M_{D}+m_{\pi}$. The dependence of the $D$- and $D^{*}$-meson masses on $m_{\pi}$ was explored in Ref. [55] using unitarized SU(3) chiral perturbation theory. The tree-level expressions for $M_{D}$ and $M_{D^{*}}$, derived from the expansion around the physical masses with the charm quark mass kept at its physical value, are as follows
@@ 3.1 LO results
The analytic structure of the scattering amplitude in the complex momentum plane ($k$-plane) is continuously changing with varying pion mass. At the nominal $D^{*}D$ threshold, the OPE has both a real and an imaginary part as long as the three-body $DD\pi$ threshold is below the $DD^{*}$ threshold. This is fulfilled, in particular, for the physical pion mass ($\xi=1$). Defining the on-shell momentum relative to the two-body threshold by $k$, with $E=M_{D}+M_{D^{*}}+k^{2}/(2\mu)$, and introducing $\Delta M=M_{D^{*}}-M_{D}$, the three-body branch point can be found by requiring $E=2M_{D}+m_{\pi}$. This leads to
@@
This can be derived from $D_{1}$ using that $\mu\approx m_{D}/2$, leading to the conclusion that $\omega(4k^{2})=\Delta M$ for backward scattering.
Figure 3: Left panel: Pion mass dependence of the $T_{cc}^{+}$ pole in the complex $k$-plane predicted at LO in chiral EFT. The value of $\xi$ is indicated by color. Right panel: LO pole trajectory in the complex energy plane as a function of $\xi$, corresponding to the left panel. Re $E_{\rm pole}$ and Im $E_{\rm pole}$ are shown by solid and dashed lines, respectively. Only the energy of the pole with $\text{Re}(k)>0$ is shown after the resonance appears.
@@
We now investigate the trajectory of the $T_{cc}^{+}$ pole as $\xi$ is varied, which is largely influenced by the non-trivial motion of the branch points described in the previous paragraph. The $T_{cc}^{+}$ pole at the physical pion mass can be interpreted as a quasi-bound state – a would be bound state of $DD^{*}$ if there were no three-body decay to $DD\pi$. When the pion mass increases, the $T_{cc}^{+}$ width decreases accordingly and the corresponding pole in the complex momentum plane ($k$-plane) approaches the imaginary axis, finally turning into a bound state. This is illustrated in Fig. 3 – see the zoomed plot in the left panel. The proper bound state occurs when the $T_{cc}^{+}$ pole, located below the $DD^{*}$ threshold, coincides with the three-body threshold. This takes place at $\xi=\xi^{\prime}\approx 1.027$, which is just a bit smaller than $\xi_{0}$. Therefore, there is a very narrow range of $\xi_{0}>\xi>\xi^{\prime}$, where the three-body threshold is still below the two-body threshold, but the $T_{cc}^{+}$ is already stable. By further departing from the physical point in terms of $\xi$, the bound state on the physical Riemann sheet (RS-I) turns into a virtual state on RS-II. The particular value of $\xi$ when this happens depends on the dynamics, namely on whether the LO or NLO potential is employed: at LO of the chiral EFT expansion, the transition emerges at $\xi\approx 1.3$. A common feature of all settings is the appearance of the second (lower-lying) virtual pole, the dynamics of which is interrelated with the location of the lhc from the OPE. The general pattern is illustrated in Fig. 3 and is as follows: At some $\xi$, the second virtual pole occurs from under the lhc branch cut and goes along with the branch point until the first (upper) pole comes close. Then, both poles collide and the state becomes a resonance. The OPE plays a significant role, not only by providing a repulsion, which would be absent in a pure contact formulation, but also affecting the analytic properties of the $DD^{*}$ scattering amplitude in a very nontrivial way. The right panel of Fig. 3 shows the behavior of the pole in the energy plane that corresponds to the $k$-plane pole trajectory in the left panel. The cusps and the point where $E_{\rm pole}=0$ indicate a change in the character of the pole, which transitions from a quasi-bound to bound, virtual and finally to a resonance state as $\xi$ increases. The point where $E_{\rm pole}=0$ corresponds to the transition from RS-I to RS-II. Qualitatively, the pole trajectories described here are similar to those discussed in Ref. [57], however, modified by the effects of the dynamical pions that induce the imaginary part of the $T_{cc}$ pole in physical setting and steer the $\xi$ dependence of the lower virtual pole.
@@ 3.2 NLO results
While the results at LO capture the behavior of the pole trajectory qualitatively, in this section, we discuss how the results change when the higher-order contact interactions in Eq. (2), $C_{2}$ and $D_{2}$, are incorporated. First, we note that the experimental data [9, 10] do not allow us to fix the momentum-dependent ${\cal O}(Q^{2})$ LEC $C_{2}$ — or stated otherwise, at the physical point there is a strong correlation between the values of $C_{2}$ and $C_{0}$. Accordingly, the EFT-based analysis of Ref. [25] resulted in an excellent description of the experimental line shapes using $C_{2}=0$. On the other hand, this momentum-dependent short-range interaction was found to be important in Refs. [25, 40] for understanding the results of Ref. [20] at $m_{\pi}=280$ MeV ($\xi\approx 2$). To extract the unknown parameters of the short-range interactions at ${\cal O}(Q^{2})$ — specifically, $C_{2}$ and $D_{2}$ in Eq. (2) — we use the phase shifts extracted in [40] from the lattice energy levels at $\xi\approx 2$ [20]. It is important to emphasize that the LECs extracted in Ref. [40] cannot be used directly in the current analysis, since Ref. [40] approximates the full $\pi DD$ Green function, as per Eq. (4), with the static pion propagator. While this approximation is justified for analyzing lattice data at $m_{\pi}=280$ MeV, the three-body cut emerging from the $\pi DD$ Green function is crucial for maintaining the correct analytic structure of the $DD^{*}$ scattering amplitude and thus for providing chiral extrapolations from the lattice data to the physical point. The results of our best fit, including the $1\sigma$ uncertainty band, are shown in Fig. 4, along with the original phase shifts from Ref. [40]. To propagate the uncertainties from the original dataset into our calculations, we use the bootstrap procedure — see, e.g., Ref. [58] for details. Specifically, we randomly generate 1000 datasets, with each data point of every data set falling within the original $1\sigma$ error range for the phase shifts from Ref. [40]. Each of these simulated datasets is individually used to determine the best-fit parameters $C_{2}$ and $D_{2}$. The resulting distribution of $\{C_{2},D_{2}\}$ from these fits is then propagated to estimate the uncertainty of the results for the $DD^{*}$ scattering amplitude and the $T_{cc}$ pole position. This uncertainty is associated with the statistical uncertainty of the finite-volume energy levels from Ref. [20].
@@
Figure 6: Comparison of the $T_{cc}$ pole trajectory at NLO in chiral EFT with that in a pionless (contact) theory (dot-dashed line). Notation is the same as in Fig. 5. Only the real part is shown, as the pole in the contact theory is always real. The arrow indicates the pion mass ($\xi\approx 1.68$), after which the pole in the pionful theory becomes a resonance state while the pole in the pionless theory remains a virtual state. The blue data point corresponds to a virtual state extracted by the Hadron Spectrum Collaboration at $m_{\pi}=391$ MeV [22].
@@
In Fig. 6, we compare our NLO results with the pole trajectory obtained from a pure contact theory using only the contact potential from Eq. (2) (see the dashed-dotted line). The LECs for the contact potential were obtained by reproducing the physical value of the $T_{cc}$ pole position and the phase shifts of Ref. [20], which were extracted from the finite-volume spectra using the Lüscher method and analyzed employing the ERE. We note that a recent lattice investigation of coupled $DD^{*}-D^{*}D^{*}$ scattering at $m_{\pi}=391$ MeV ($\xi\approx 2.85$) [22], also using the Lüscher method and ignoring the lhc, predicts a virtual state $62\pm 34$ MeV below the $DD^{*}$ threshold, fully consistent with the about 80 MeV virtual state predicted by our dot-dashed curve (please note that we do not provide uncertainty for the pionless case). On the other hand, it is evident that the additional repulsion from the longest-range OPE potential significantly impacts the results. Unlike the contact trajectory, which remains a virtual state with growing pion mass, our NLO pole position transforms from a virtual state to a resonance at pion masses corresponding to $\xi\approx 1.7$.

### v2

@@ 1 Introduction
In 2021, the LHCb experiment discovered the first exotic doubly-charmed narrow resonance, denoted as $T_{cc}(3875)^{+}$, whose minimal quark composition is $cc\bar{u}\bar{d}$ [9, 10]. This discovery revealed a state with a mass just a few hundred keV below the $D^{*+}D^{0}$ threshold, with a width primarily dictated by its strong decay mode to $DD\pi$. Since approximately 90% of the $D^{0}D^{0}\pi^{0}$ events contain a genuine $D^{*}$ meson [10], it is natural to expect (see the discussions in Refs. [11, 12, 13]) that the width of the $T_{cc}(3875)^{+}$ should be smaller than that of the $D^{*}$, which is only $(83.4\pm 1.8)$ keV [14]. The properties of this state have been investigated employing low-energy effective field theories (EFT) [15, 16, 17, 18, 19, 20] and phenomenological models, see, e.g., [7] and references therein. In particular, the pole position of the $T_{cc}(3875)^{+}$ and the $DD^{*}$ scattering parameters were extracted in Ref. [17] using the leading order chiral EFT approximation by performing a coupled-channel analysis of the experimental line shape in the $D^{0}D^{0}\pi^{+}$ final state. In that work, a special attention was paid to the inclusion of three-body cuts, which were found to be crucial for the accurate determination of the $T_{cc}$ pole position in the complex energy plane. We emphasize that for the $T_{cc}$, including the one-pion exchange (OPE) to all orders with a proper treatment of its cuts is needed for theoretical consistency. To illustrate this, we note that the $D^{*}$ self-energy induces an imaginary part to the $D^{*}D$ propagator, which corresponds to the $DD\pi$ intermediate state going on shell. This on-shell intermediate state contains a pair of identical mesons ($DD$) in isospin 1, which has to be in an even partial wave to comply with Bose symmetry. Proper symmetrization of the $DD$ state, that ensures this, generates a one-pion exchange. Moreover, since the $D^{*}$ self-energy is resummed in the $D^{*}D$ propagator, consistency with Bose symmetry requires resumming the OPE as well. For a similar discussion on the Pauli principle in the context of two identical nucleons in the three-body intermediate state, see Refs. [21, 22, 23].
@@
In this work, we employ the same chiral EFT-based approach to scrutinize the analytic structure of the scattering amplitude as a function of the pion mass. Given the intricate interplay between the right-hand (three-body) and left-hand cuts, which depend sensitively upon the pion mass, we incorporate all energy and/or momentum scales relevant for this complicated dynamics to predict the pole trajectory of the $T_{cc}$ state for pion masses between the physical value $m_{\pi}^{\rm ph}$ and $3m_{\pi}^{\rm ph}$. This information can serve as a benchmark for future lattice QCD calculations and is important for obtaining additional insights into the structure of $T_{cc}$ state [47]. Our analysis also yields predictions for the $DD^{*}$ phase shifts at any value of $m_{\pi}$ within the considered range, in spite of the fact that the ERE has a very limited range of validity. The lhc also necessitates an improvement of the Weinberg approach to properly describe the compositeness of a hadronic state. Besides the fact that the zeros in the $T$-matrix emerging from the interplay of the repulsive OPE potential and attractive short-range physics may invalidate the original Weinberg formalism, as discussed in Refs. [48, 49, 50], already the small scale introduced into the system by the nearby lhc calls for a refined formulation of the compositeness criterion.
@@ 2 Framework
$V_{\text{OPE}}(E,{\bm{p}},{\bm{p}}^{\prime})=-\frac{g^{2}}{8f_{\pi}^{2}}\frac{({\bm{q}\cdot\bm{\epsilon}})({\bm{q}\cdot\bm{\epsilon}^{\prime*}})}{2\omega_{\pi}({\bm{q}}^{2})}\,D^{\pi}(E,{\bm{p}},{\bm{p}}^{\prime}),$ (3)
@@
The partial-wave-projected potentials $V_{\alpha\beta}(E,p,p^{\prime})$ are obtained along the lines of Refs. [57, 58] as follows,
@@
$+\int_{0}^{\Lambda}\frac{\text{d}{q}\,q^{2}}{2\pi^{2}}V_{\alpha\gamma}(E,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p},q)G(E,q)T_{\gamma\beta}(E,q,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p^{\prime}}).$
@@
$\Gamma(E,q)=\frac{g^{2}M_{D}}{8\pi f_{\pi}^{2}M_{D^{*}}}\Big{[}\Sigma(s)-\Sigma_{0}(s)\theta(M_{D}+m_{\pi}-M_{D^{*}})\Big{]}$
is the dynamical width of the $D^{*}$ with
@@
and $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2bc-2ca$ is the Källén triangle function, and $s=[E-M_{D}-q^{2}/(2\mu)]^{2}$. Here
@@
where the first and second terms renormalize the $D^{*}$ mass and wave function, respectively, if $M_{D^{*}}<M_{D}+m_{\pi}$. The dependence of the $D$- and $D^{*}$-meson masses on $m_{\pi}$ was explored in Ref. [59] using unitarized SU(3) chiral perturbation theory. The tree-level expressions for $M_{D}$ and $M_{D^{*}}$, derived from the expansion around the physical masses with the charm quark mass kept at its physical value, are as follows,
@@ 3.1 LO results
The analytic structure of the scattering amplitude in the complex momentum plane ($k$-plane) is continuously changing with varying pion mass. At the nominal $D^{*}D$ threshold, the OPE has both a real and an imaginary part as long as the three-body $DD\pi$ threshold is below the $DD^{*}$ threshold. This is fulfilled, in particular, for the physical pion mass ($\xi=1$). Defining the on-shell momentum relative to the two-body threshold by $k$, with $E=M_{D}+M_{D^{*}}+k^{2}/(2\mu)$, and introducing $\Delta M=M_{D^{*}}-M_{D}$, the three-body branch point can be found by requiring $E_{\rm rhc_{3}}{=M_{D}+M_{D^{*}}+k^{2}_{\rm rhc_{3}}/(2\mu)}=2M_{D}+m_{\pi}$. This leads to
@@
This can be derived from $D_{1}$ using that $\mu\approx m_{D}/2$, leading to the conclusion that $\omega(4k^{2})=\Delta M$ for forward $DD^{*}$ scattering ($\cos{({\bm{p}}\cdot{\bm{p}}^{\prime})}=1$).
Figure 3: Left panel: Pion mass dependence of the $T_{cc}$ pole in the complex $k$-plane predicted at LO in chiral EFT. The value of $\xi$ is indicated by color. The second, more distant virtual state is also shown, along with its collision point with the $T_{cc}$ pole, after which the resonance poles emerge. Right panel: LO trajectory of the $T_{cc}$ pole in the complex energy plane as a function of $\xi$, corresponding to the left panel. Re $E_{\rm pole}$ and Im $E_{\rm pole}$ are shown by solid and dashed lines, respectively. The second, more distant state is not shown; only the energy of the pole with $\text{Re}(k)>0$ is shown after the resonance appears.
@@
We now investigate the trajectory of the $T_{cc}^{+}$ pole as $\xi$ is varied, which is largely influenced by the non-trivial motion of the branch points described in the previous paragraph. The $T_{cc}^{+}$ pole at the physical pion mass can be interpreted as a quasi-bound state – a would be bound state of $DD^{*}$ if there were no three-body decay to $DD\pi$. When the pion mass increases, the $T_{cc}^{+}$ width decreases accordingly and the corresponding pole in the complex momentum plane ($k$-plane) approaches the imaginary axis, finally turning into a bound state. This is illustrated in Fig. 3 – see the zoomed plot in the left panel. Please note, however, that the imaginary part of the quasi-bound state energy is so small that it is indistinguishable from zero in the right panel of Fig. 3. The proper bound state occurs when the $T_{cc}^{+}$ pole, located below the $DD^{*}$ threshold, coincides with the three-body threshold. This takes place at $\xi=\xi^{\prime}\approx 1.027$, which is just a bit smaller than $\xi_{0}$. Therefore, there is a very narrow range of $\xi_{0}>\xi>\xi^{\prime}$, where the three-body threshold is still below the two-body threshold, but the $T_{cc}^{+}$ is already stable. By further departing from the physical point in terms of $\xi$, the bound state on the physical Riemann sheet (RS-I) turns into a virtual state on RS-II. The particular value of $\xi$ when this happens depends on the dynamics, namely on whether the LO or NLO potential is employed: at LO of the chiral EFT expansion, the transition emerges at $\xi\approx 1.3$. A common feature of all settings is the appearance of the second (lower-lying) virtual pole, the dynamics of which is interrelated with the location of the lhc from the OPE. The general pattern is illustrated in Fig. 3 and is as follows: At some $\xi$, the second virtual pole occurs from under the lhc branch cut and goes along with the branch point until the first (upper) pole comes close. Then, both poles collide and the state becomes a resonance. The OPE plays a significant role, not only by providing a repulsion, which would be absent in a pure contact formulation, but also affecting the analytic properties of the $DD^{*}$ scattering amplitude in a very nontrivial way. The right panel of Fig. 3 shows the behavior of the pole in the energy plane that corresponds to the $k$-plane pole trajectory in the left panel. The cusps and the point where $E_{\rm pole}=0$ indicate a change in the character of the pole, which transitions from a quasi-bound to bound, virtual and finally to a resonance state as $\xi$ increases. The point where $E_{\rm pole}=0$ corresponds to the transition from RS-I to RS-II. Qualitatively, the pole trajectories described here are similar to those discussed, e.g., in Refs. [61, 62, 63, 64, 47], though modified by the effects of dynamical pions. These pionic effects induce the only hadronic contribution to the imaginary part of the $T_{cc}$ pole at the physical point and influence the $\xi$ dependence of the $T_{cc}$ pole as well as the second more distant virtual pole.
@@ 3.2 NLO results
While the results at LO capture the behavior of the pole trajectory qualitatively, in this section, we discuss how the results change when the higher-order contact interactions in Eq. (2), $C_{2}$ and $D_{2}$, are incorporated. First, we note that the experimental data [9, 10] do not allow us to fix the momentum-dependent ${\cal O}(Q^{2})$ LEC $C_{2}$ — or stated otherwise, at the physical point there is a strong correlation between the values of $C_{2}$ and $C_{0}$. Accordingly, the EFT-based analysis of Ref. [29] resulted in an excellent description of the experimental line shapes using $C_{2}=0$. On the other hand, this momentum-dependent short-range interaction was found to be important in Refs. [29, 44] for understanding the results of Ref. [24] at $m_{\pi}=280$ MeV ($\xi\approx 2$). To extract the unknown parameters of the short-range interactions at ${\cal O}(Q^{2})$ — specifically, $C_{2}$ and $D_{2}$ in Eq. (2) — we use the phase shifts extracted in [44] from the lattice energy levels at $\xi\approx 2$ [24]. It is important to emphasize that the LECs extracted in Ref. [44] cannot be used directly in the current analysis, since Ref. [44] approximates the full $\pi DD$ Green function, as per Eq. (4), with the static pion propagator. While this approximation is justified for analyzing lattice data at $m_{\pi}=280$ MeV, the three-body cut emerging from the $\pi DD$ Green function is crucial for maintaining the correct analytic structure of the $DD^{*}$ scattering amplitude and thus for providing chiral extrapolations from the lattice data to the physical point. Thus, we adjust the contact terms to the phase shifts with the main criterion that the resulting central curve and its uncertainty band should resemble those from Ref. [44]. The results of our best fit, including the $1\sigma$ uncertainty band, are shown in Fig. 4, along with the original phase shifts from Ref. [44], and are in good agreement. However, we note that we do not account for possible correlations between the input data points. This leads to a more conservative estimate of the propagated uncertainty. To propagate the uncertainties from the original dataset into our calculations, we use the bootstrap procedure — see, e.g., Ref. [65] for details. Specifically, we employ the orange band in Fig. 4 from Ref. [44] to randomly generate 1000 datasets, assuming a Gaussian distribution. Each of these simulated datasets is individually used to determine the best-fit parameters $C_{2}$ and $D_{2}$. The resulting distribution of $\{C_{2},D_{2}\}$ from these fits is then propagated to estimate the uncertainty of the results for the $DD^{*}$ scattering amplitude and the $T_{cc}$ pole position. This uncertainty is associated with the statistical uncertainty of the finite-volume energy levels from Ref. [24].
@@
Figure 6: Comparison of the $T_{cc}$ pole trajectory at NLO in chiral EFT with that in a pionless (contact) theory (dot-dashed line). Notation is the same as in Fig. 5. Only the real part is shown, as the pole in the contact theory is always real; gray band indicates the statistical uncertainty of the contact EFT calculation propagated from fits to the lattice energy levels based on the ERE [24]. The arrow indicates the pion mass ($\xi\approx 1.68$), after which the pole in the pionful theory becomes a resonance state while the pole in the pionless theory remains a virtual state. The blue data point corresponds to a virtual state extracted by the Hadron Spectrum Collaboration at $m_{\pi}=391$ MeV [26], while the green square corresponds to a virtual state at $m_{\pi}\approx 348.5$ MeV, extracted using the ERE parameters from Ref. [25].
@@
In Fig. 6, we compare our NLO results with the pole trajectory obtained from a pure contact theory using only the contact potential from Eq. (2) (see the dashed-dotted line). The LECs for the contact potential were obtained by reproducing the physical value of the $T_{cc}$ pole position and the phase shifts of Ref. [24], which were extracted from the finite-volume spectra using the Lüscher method and analyzed employing the ERE. We note that a recent lattice investigation of coupled $DD^{*}-D^{*}D^{*}$ scattering at $m_{\pi}=391$ MeV ($\xi\approx 2.85$) [26], also using the Lüscher method and ignoring the lhc, predicts a virtual state $62\pm 34$ MeV below the $DD^{*}$ threshold, fully consistent with the about 75 MeV virtual state predicted by our dot-dashed curve. In addition, the virtual state pole obtained by extrapolating the ERE from Ref. [25] to the below-threshold region (see the green point) is also in line with the prediction from our contact EFT. On the other hand, it is evident that the additional repulsion from the longest-range OPE potential, not included in Refs. [26, 25], significantly impacts the results. Unlike the contact trajectory, which remains a virtual state with growing pion mass, our NLO pole position transforms from a virtual state to a resonance at pion masses corresponding to $\xi\approx 1.7$.